OpenStudy (anonymous):

Does anyone know how to compute these following combinations? 1. C(6,4) 2. C(7,5) 3. C(9,6) 4. C(8,2) x C(8,6)

OpenStudy (campbell_st):

the formula is \[nCr = (n!)/((r!)\times(n-r)!)\] \[6C4 = 6!/((4!)\times(6-4)!)\] You should have a calculator function that does it for you....

OpenStudy (anonymous):

yea but we cant use calculators for the test

OpenStudy (campbell_st):

its factorial notation (6x5x4x3x2x1)/((4x3x2x1)x(2x1)) cancel common factors and you'll be left with 3 x 5 = 15

OpenStudy (anonymous):

why the 2x1 at the end of the equation come from

OpenStudy (campbell_st):

because the formula says (n-r)! this is 6 - 4)! or (2)!

OpenStudy (anonymous):

ok thank you very much